OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (jim766):

x^2 - 36 = 0 you can solve this one by getting the x^2 by itself +36 + 36 x^2 = 36 now take the sq rt of both sides

OpenStudy (jim766):

what number times itself = 36

OpenStudy (whpalmer4):

For \(x^2=8x-12\) or \(x^2-8x+12 = 0\) you can solve with the quadratic equation, by factoring, or by completing the square. Factoring is probably the easiest.

OpenStudy (anonymous):

How do you describe the solution and how are the alike/different?

OpenStudy (whpalmer4):

For the first one, be aware that there are two answers, not one, because there are two numbers which when multiplied by themselves give 36...

OpenStudy (anonymous):

6 is one of them

OpenStudy (whpalmer4):

Yes. And the other is very similar :-)

OpenStudy (whpalmer4):

What do you get if you square a negative number?

OpenStudy (anonymous):

-6?

OpenStudy (whpalmer4):

Right! -6 is the other solution to \(x^2=36\)

OpenStudy (anonymous):

so how would I write that?

OpenStudy (whpalmer4):

\(x = \pm 6\) would be one way. Or \(x=6, x=-6\)

OpenStudy (anonymous):

Math is not my subject I don't understand it and its killing me. I have to get my online math class finished by sunday

OpenStudy (whpalmer4):

Okay, let's solve the other one, because I have to leave shortly. Do you know how to do any of the 3 methods I mentioned?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

I still dont understand how to write them out as solved and part 2 or 3

OpenStudy (whpalmer4):

Okay, let's do the quadratic, then, because it always works, even if it might be more work sometimes. If you have a quadratic equation \(ax^2+bx+c=0\), with \(a\ne0\), then the solutions are given by the quadratic formula:\[x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]and it is just a matter of plugging in the numbers and doing the arithmetic. For that first equation, which we've already solved, we had \(x^2 - 36 = 0\) which we can rewrite as \(1x^2 + 0x -36 = 0\) and so \(a = 1, b = 0, c = -36\) Plugging into the formula, that gives us\[x = \frac{-(0)\pm\sqrt{(0)^2-4(1)(-36)}}{2(1)}\]and that simplifies to \[x=\frac{\pm\sqrt{144}}{2} = \frac{\pm12}{2} = \pm 6\]

OpenStudy (whpalmer4):

For the second equation, we have to rearrange it a little bit to put it in that form: \[x^2 = 8x – 12\]Subtract 8x from both sides\[x^2-8x=\cancel{8x}-\cancel{8x}-12\]Add 12 to both sides\[x^2-8x+12=0\]Now we see that \(a = 1, b=-8, c = 12\) and I'll let you plug them into the formula.

OpenStudy (anonymous):

how are they alike /different?

OpenStudy (whpalmer4):

What do the solutions represent? Well, a picture may be worth a few words here: This is the graph of \(y = x^2-36\)

OpenStudy (anonymous):

it represents 6, -6

OpenStudy (whpalmer4):

You should make a graph of the other equation (\y = x^2-8x+12\) for comparison.

OpenStudy (whpalmer4):

You're on the right track — the point where the curve crosses the x-axis are the solutions to the equation.

OpenStudy (whpalmer4):

the points, in this case

OpenStudy (anonymous):

-33?

OpenStudy (whpalmer4):

No, -33 isn't a solution to either of those equations.

OpenStudy (whpalmer4):

How did you get that?

OpenStudy (anonymous):

on the picture at the bottom it curves at -33 on the y axis

OpenStudy (whpalmer4):

x-axis, not y-axis, and those ticks are 2, not 1 (count the ticks between 0 and -10)

OpenStudy (anonymous):

okay so nevermind on that one.

OpenStudy (whpalmer4):

You were asked to solve \(x^2-36=0\). What I graphed was \(y = x^2-36\). At the points where \(y = 0\), we are at a value of \(x\) such that \(y = 0 = x^2-36\). Therefore, those points are the solutions to \(x^2-36=0\). Does that make sense?

OpenStudy (whpalmer4):

Why don't you try solving the other equation so I can check your work? I have to leave in a few minutes.

OpenStudy (anonymous):

where do I use a b and c?

OpenStudy (whpalmer4):

In the formula I gave you...

OpenStudy (anonymous):

For the second equation, we have to rearrange it a little bit to put it in that form: x2=8x–12 Subtract 8x from both sides x2−8x=8x−8x−12 Add 12 to both sides x2−8x+12=0 Now we see that a=1,b=−8,c=12 and I'll let you plug them into the formula.

OpenStudy (anonymous):

this is the formula

OpenStudy (anonymous):

add 12 to to 12 and 0?

OpenStudy (anonymous):

so 12 and 24?

OpenStudy (whpalmer4):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] sorry, gotta go!